two gallons of milk contains 2% butterfat. How many quarts of milk should be drained and replaced by pure butterfat to produce a mixture of 2 gallons of milk containing 5% butterfat?
Answer by jorel1380(3719)

2x.02=.04
2x.05=.1
m+n=2
.02m+n=.1
2m+100n=10
2m+2n=4
98n=6
n=6/98 gallons=.06122
(1.9387755x.02)+.06122=.1..